# Generalized Compressed Sensing for Image Reconstruction with Diffusion Probabilistic Models

### v3

@@ 3.4 Characterizing the optimized linear measurements
Figure 4: Optimized measurement vectors are distinct from PC and random measurements. A) Left: Twenty-five selected principal components of the MNIST dataset, visualized as images, sorted by variance explained from top to bottom, and left to right. Right: Twenty-five components from the columns of the OLM for $k=64$. B) Same as A) but for the CelebA dataset. C) The variance of the measurement on each component as a function of number of measurements ($k$) for the PCs (blue) and OLMs (red). Note that the optimal measurements are sorted post hoc based on the variance explained by each component. These vectors are optimized jointly for each $k$ and are not ordered by the algorithm itself. D) The Grassmann distance between subspaces spanned by different linear measurements as a function of $k$ (Hamm & Lee, 2008). We show the distance between subspaces spanned by PCs obtained on two random halves of the training data (blue), between subspaces spanned by PCs and OLMs (red), and between subspaces spanned by PCs and a set of random measurements (green). See Appendix E for the definition of Grassmann distance. E) The (absolute) skewness value of the measurement on each component as a function of number of measurements ($k$, sorted post hoc based on the skewness value) for the PCs (blue) and OLMs (red). F) Distribution of mutual information between all pairs of measurement directions for PCA (blue), ICA (green), and OLM (red). G) Performance (on MNIST) and visualization of OLMs (olive, $\text{OLM}_{\text{seq}}$) obtained via a sequential version of the algorithm Section E.6. H) Performance as a function of measurement noise (on MNIST, $k=32$) and visualization of OLMs obtained with Gaussian noise ($\sigma=0.4$) added to the linear measurements during optimization (olive, $\text{OLM}_{\text{noise}}$), along with comparison to the regular OLMs (red) and PCs (orange).
@@
In this section, we present some qualitative and quantitative analyses to help interpret the difference between the linear measurement subspaces defined by PCs and OLMs. For the MNIST dataset, as expected, the first few PCs appear digit-like, and the measurement vectors contain increasing high-spatial frequency content as $k$ increases (Figure 4A, left). The OLMs, on the other hand, do not follow this coarse-to-fine (low to high frequency) ordering and are visually much more similar to each other (Figure 4A, right). This is reflected, quantitatively, in the measurement variance as a function of $k$, which falls rapidly for the PCs, but slowly for the OLMs (Figure 4C, left).
@@
Similar phenomena are observed for the CelebA dataset (Figure 4B, C). The PCs, known as “eigenfaces” (Sirovich & Kirby, 1987; Turk & Pentland, 1991), contain features that are geometrically aligned with features of the face, with increasing spatial frequency content. The variance of the measurements across PCs falls approximately exponentially. Although the OLMs also have a face-like appearance, they differ from the PCs. Each OLM vector has a detailed representation of fine facial features (eyes, nose, lips), and a coarse representation of the surrounding content (forehead, chin, hair). As in the MNIST case, all OLMs have similar spatial frequency composition. Consistent with this, the variance of the measurements falls slowly across the OLMs (Figure 4C, right).
@@
To quantify the differences between these two sets of measurement vectors, we use the Grassmann distance between the subspaces they span (Hamm & Lee, 2008; Björck & Golub, 1973). Figure 4D shows this distance as a function of $k$. As a control, the blue line measures the distance between subspaces arising from PCs computed on two random halves of the training data. As expected, the distance between these two subspaces is small. On the other hand, for both the MNIST and CelebA dataset, the subspaces spanned by the optimal measurements are distinct from those defined by the PCs (Figure 4D, red), but they are more similar to the PCs than to the space defined by a set of random measurements (Figure 4D, green).
@@
We can also characterize the linear measurements by analyzing the distribution of measurement values they produce in response to natural images (i.e., the measurement distribution). We found that OLMs differ from traditional methods in two key aspects: First, their measurement distributions are highly asymmetrical, as indicated by the significant non-zero skewness values (see Figure 4E). Second, unlike PCA or ICA, OLMs produce non-factorized measurement distributions: the measurements across axes are not fully independent. This dependency is reflected in the increased pairwise mutual information between different measurement directions observed for OLMs compared to those obtained with PCA or ICA (Figure 4F). This observation aligns with the intuition from the 2D example in Figure 1, where optimal measurements for non-Gaussian priors are not necessarily aligned with the axes.
@@
Lastly, we also consider two modifications to the optimization procedure that result in different OLMs. First, in Figure 4G we present a set of OLMs that are sequentially organized – similar to PCs – using a greedy version of the algorithm (see Section E.6), where the measurement subspace for smaller $k$ is always contained within that of larger $k$. We found that while enforcing this sequential constraint results in a slight decrease in performance, the resulting OLMs exhibit more ordered spatial frequency content. Nonetheless, they remain distinct from the PCs both visually and when comparing subspace distances (not shown).
@@
Second, in Figure 4H we examine the robustness of different measurement matrices to noise. For both PCs and OLMs, PSNR decreases as an increasing amount of Gaussian noise is added to the measurements. The OLMs can be made more robust to noise by injecting measurement noise during the OLM optimization (in this case, with a fixed standard deviation of $\sigma=0.4$). This results in a more noise-robust set of linear measurements ($\text{OLM}_{\text{noise}}$), accompanied by a modest performance loss in low-noise regimes. The noise-optimized OLMs are less smooth than those optimized without noise (Figure 4H). A more thorough investigation of noise robustness is worthwhile, but will require the development of a reconstruction algorithm that operates on noisy measurements.
@@ 3.5 Optimized linear measurements for perceptual loss
As an example, we computed a set of OLMs with respect to structural similarity index (SSIM) (Wang et al., 2004), a widely used perceptual image quality metric. Figure 5 shows results obtained using this set of optimal linear measurements. We obtained a noticeable improvement in mean SSIM for a range of $k$, and this improvement applies to the majority of the images in the test set. In Figure 5B, four example images are shown to highlight the visual improvement obtained by optimizing for the SSIM objective. However, it is worth noting that optimizing for SSIM does result in a lower average PSNR compared to the PCs, suggesting a trade-off between optimizing for perceptual quality versus MSE reconstruction, at least with our methods.
Lastly, in Figure 5C, the SSIM-optimized measurement vectors are visualized, which appear quite distinct from both the PCs and their MSE-optimized counterpart shown in Figure 4B. We further confirmed this difference by analyzing the measurement subspace, and found that the subspace distance between OLMs optimized toward different objectives is comparable to their distance from the PCs.
@@ 4 Related Work
The main advantages of our method are that (1) the linear measurements are optimized directly with respect to the end objective of the reconstruction problem, and that (2) we incorporate a highly expressive learned prior to exploit the higher-order statistics of natural images. As was shown in Section 2, our method outperforms both random measurements and PCA.
@@ Appendix A Tweedie / Miyasawa Proof
For completeness, we provide the proof of the core relationship underlying diffusion models (Miyasawa, 1961; Robbins, 1992; Efron, 2011; Raphan & Simoncelli, 2007). Consider the problem of estimating a natural image $x$ given a noise-corrupted measurement $y=x+\epsilon$, with $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$. The estimate $\hat{x}(y)$ is chosen to minimize the mean of the squared error $||x-\hat{x}(y)||_{2}^{2}$ over all images $x$ and observations $y$ corrupted with known $\sigma^{2}$. In Bayesian terms, the solution to this denoising problem is the posterior mean:
[…]
@@
Thus, the residual of the optimal denoiser, $\hat{x}(y)-y$ is proportional to the gradient of $\log p(y)$. Note that $p(y)$ is equal to the signal prior $p(x)$ convolved with (blurred by) a Gaussian with variance $\sigma^{2}$. This quantity can be used to sample from the implicit prior $p(x)$, for example, by using annealed Langevin dynamics, where the noise $\sigma^{2}$ corresponds to different temperature levels Song & Ermon (2019); Bussi & Parrinello (2007); Bhattacharya & Waymire (2009). See Appendix D for a detailed description of our sampling procedure.
@@ Appendix B Principal Components Analysis
Denote by $X$ the data matrix where the columns are individual mean-subtracted samples $x_{i}$. The linear projection is defined as $\hat{x_{i}}=MM^{T}x_{i}$, where $M$ is a matrix with orthogonal columns. Plugging this expression into Equation 5, we have the expected squared-error loss for linear projection:
$l(M)$ $=\frac{1}{N}\sum_{i}||x_{i}-MM^{T}x_{i}||_{2}^{2}$ (14)
@@
When the data is in fact Gaussian distributed, linear projection onto the first $k$ PCs does achieve the minimal error possible, as shown in the bivariate example in Fig. 1A. For non-Gaussian distributions, nonlinear estimation can potentially improve performance.
@@ Appendix C Sampling Algorithm
6:     $s_{t}\leftarrow f(x_{t-1})-x_{t-1}$triangleright Compute the score from the denoiser residual
7:     $\sigma_{t}^{2}\leftarrow||s_{t}||^{2}/{d}$triangleright Compute the current noise level for stopping criterion
@@
10:     $x_{t}\leftarrow x_{t-1}+hd_{t}+\gamma_{t}z_{t}$triangleright Perform a partial denoiser step and add noise
@@ Appendix D Constrained Sampling Algorithm
2:$t\leftarrow 1$triangleright initialization
3:$\mu\leftarrow 0.5(\mathbf{1}-MM^{T}\mathbf{1})+Mm$
4:$y_{1}\leftarrow\mathcal{N}(\mu,\ I)$
5:$\sigma^{2}_{1}\leftarrow||g(y_{1})||^{2}_{2}/d$
6:while $\ \sigma_{t}>\sigma_{end}$ do $l_{t}\leftarrow(I-MM^{T})g(y_{t})+M(m-M^{T}y_{t})$triangleright compute the conditional gradient $\gamma^{2}\leftarrow[{(1-\beta*h)^{2}-(1-h)^{2}}]*\sigma^{2}_{t}$triangleright scale factor for the added noise $y_{t+1}\leftarrow y_{t}+h*l_{t}+\gamma*\mathcal{N}(0,I)$triangleright move up the gradient and add noise $\sigma^{2}_{t+1}\leftarrow||l_{t}||^{2}_{2}/d$triangleright compute an estimated noise level $t\leftarrow t+1$
@@ E.3 MMSE estimate
We construct the linear inverse estimate by averaging multiple samples obtained using Algorithm 2. The individual samples are shown in Figure 2 and Supplementary Figure 3 and Figure 4. Averaging multiple samples achieves a lower MSE by approximating the posterior mean, which we use when reporting performance (PSNR values). The number of samples $n$ is set to $n=2$ when optimizing the linear measurement to optimize performance, as we found empirically that using more samples for the MMSE estimates during optimization does not improve performance during testing. For reporting performance on test data, we used $n=16$.
@@ E.4 Optimizing linear measurement
We search for the optimal linear measurement by performing stochastic gradient descent in the space of orthogonal matrices as described in the main text. The linear inverse procedure Algorithm 2 is end-to-end differentiable, and thus optimization of the measurement matrix can be done directly in PyTorch by taking derivatives of reconstruction loss with respect to the matrix parameterization. In all cases, we used the Adam optimizer (Kingma & Ba, 2014) with a learning rate of $10^{-4}$ with exponential decay of $0.90$. We used a batch size of $64$, and training was run for $16$ epochs. The optimization was done on a single node in a GPU cluster, with 4 NVIDIA A100 GPUs. The optimization for a single OLM requires a few hours for the MNIST dataset, and about 24 hours for the CelebA dataset. The computational cost is primarily due to backpropagation through the diffusion model sampling procedure, which involves iterative applications of the U-Net denoiser.
@@ E.5 Subspace distance
$\cos\theta_{k}=\max_{v_{k}\in F}\max_{v_{k}\in G}u^{T}_{k}v_{k},\ \ ||u_{k}||_{2}=1,||v_{k}||_{2}=1$ (15)
$\text{subject to}\ \ u^{T}_{k}u_{j}=0,v^{T}_{k}v_{j}=0,\ \forall j=1,2,...,k-1.$
@@ E.6 Sequential OLM
The algorithm presented in Section 2.4 learns an OLM independently for each value of $k$. Empirically, we observe that the linear measurement subspace of smaller $k$ is typically not fully contained in that of larger $k$. In this section, we modify the algorithm into a greedy version, where measurement vectors are added one at a time, each conditioned on the previous measurement matrix. This results in a sequentially ordered set of OLMs, analogous to the structure of PCA.
The algorithm works as follows: Given a measurement matrix $M_{k}\in R^{d\times k}$, we find a set of vectors $U=\{u_{1},u_{2},...,u_{d-k}\}$ as a basis of the null space of $M_{k}$. The new measurement vector $m^{*}$ is parameterized as $m^{*}=U\lambda$, where $\lambda$ is a unit vector. The next measurement matrix $M^{*}\in R^{d\times(k+1)}$ is obtained by concatenating $M_{k}$ with $m^{*}$ as $M^{*}=[M_{k}\ m^{*}]$. The objective function in Equation 8 can be written with respect to $\lambda$ as:
$\mathcal{L}(\lambda)=\frac{1}{N}\sum_{i=1}^{N}||h(M^{*}(\lambda)^{T}x_{i};\ M^{*}(\lambda))-x_{i}||_{2}^{2}.$ (17)

### v4

@@ 3.4 Characterizing the optimized linear measurements
Figure 4: Optimized measurement vectors are distinct from PC and random measurements. A) Left: Twenty-five selected principal components of the MNIST dataset, visualized as images, sorted by variance explained from top to bottom, and left to right. Right: Twenty-five components from the columns of the OLM for $k=64$. B) Same as A) but for the CelebA dataset. C) The variance of the measurement on each component as a function of number of measurements ($k$) for the PCs (blue) and OLMs (red). Note that the optimal measurements are sorted post hoc based on the variance explained by each component. These vectors are optimized jointly for each $k$ and are not ordered by the algorithm itself. D) The Grassmann distance between subspaces spanned by different linear measurements as a function of $k$ (Hamm & Lee, 2008). We show the distance between subspaces spanned by PCs obtained on two random halves of the training data (blue), between subspaces spanned by PCs and OLMs (red), and between subspaces spanned by PCs and a set of random measurements (green). See Appendix˜E for the definition of Grassmann distance. E) The (absolute) skewness value of the measurement on each component as a function of number of measurements ($k$, sorted post hoc based on the skewness value) for the PCs (blue) and OLMs (red). F) Distribution of mutual information between all pairs of measurement directions for PCA (blue), ICA (green), and OLM (red). G) Performance (on MNIST) and visualization of OLMs (olive, $\text{OLM}_{\text{seq}}$) obtained via a sequential version of the algorithm Section˜E.6. H) Performance as a function of measurement noise (on MNIST, $k=32$) and visualization of OLMs obtained with Gaussian noise ($\sigma=0.4$) added to the linear measurements during optimization (olive, $\text{OLM}_{\text{noise}}$), along with comparison to the regular OLMs (red) and PCs (orange).
@@
In this section, we present some qualitative and quantitative analyses to help interpret the difference between the linear measurement subspaces defined by PCs and OLMs. For the MNIST dataset, as expected, the first few PCs appear digit-like, and the measurement vectors contain increasing high-spatial frequency content as $k$ increases (Figure˜4A, left). The OLMs, on the other hand, do not follow this coarse-to-fine (low to high frequency) ordering and are visually much more similar to each other (Figure˜4A, right). This is reflected, quantitatively, in the measurement variance as a function of $k$, which falls rapidly for the PCs, but slowly for the OLMs (Figure˜4C, left).
@@
Similar phenomena are observed for the CelebA dataset (Figure˜4B, C). The PCs, known as “eigenfaces” (Sirovich & Kirby, 1987; Turk & Pentland, 1991), contain features that are geometrically aligned with features of the face, with increasing spatial frequency content. The variance of the measurements across PCs falls approximately exponentially. Although the OLMs also have a face-like appearance, they differ from the PCs. Each OLM vector has a detailed representation of fine facial features (eyes, nose, lips), and a coarse representation of the surrounding content (forehead, chin, hair). As in the MNIST case, all OLMs have similar spatial frequency composition. Consistent with this, the variance of the measurements falls slowly across the OLMs (Figure˜4C, right).
@@
To quantify the differences between these two sets of measurement vectors, we use the Grassmann distance between the subspaces they span (Hamm & Lee, 2008; Björck & Golub, 1973). Figure˜4D shows this distance as a function of $k$. As a control, the blue line measures the distance between subspaces arising from PCs computed on two random halves of the training data. As expected, the distance between these two subspaces is small. On the other hand, for both the MNIST and CelebA dataset, the subspaces spanned by the optimal measurements are distinct from those defined by the PCs (Figure˜4D, red), but they are more similar to the PCs than to the space defined by a set of random measurements (Figure˜4D, green).
@@
We can also characterize the linear measurements by analyzing the distribution of measurement values they produce in response to natural images (i.e., the measurement distribution). We found that OLMs differ from traditional methods in two key aspects: First, their measurement distributions are highly asymmetrical, as indicated by the significant non-zero skewness values (see Figure˜4E). Second, unlike PCA or ICA, OLMs produce non-factorized measurement distributions: the measurements across axes are not fully independent. This dependency is reflected in the increased pairwise mutual information between different measurement directions observed for OLMs compared to those obtained with PCA or ICA (Figure˜4F). This observation aligns with the intuition from the 2D example in Figure˜1, where optimal measurements for non-Gaussian priors are not necessarily aligned with the axes.
@@
Lastly, we also consider two modifications to the optimization procedure that result in different OLMs. First, in Figure˜4G we present a set of OLMs that are sequentially organized – similar to PCs – using a greedy version of the algorithm (see Section˜E.6), where the measurement subspace for smaller $k$ is always contained within that of larger $k$. We found that while enforcing this sequential constraint results in a slight decrease in performance, the resulting OLMs exhibit more ordered spatial frequency content. Nonetheless, they remain distinct from the PCs both visually and when comparing subspace distances (not shown).
@@
Second, in Figure˜4H we examine the robustness of different measurement matrices to noise. For both PCs and OLMs, PSNR decreases as an increasing amount of Gaussian noise is added to the measurements. The OLMs can be made more robust to noise by injecting measurement noise during the OLM optimization (in this case, with a fixed standard deviation of $\sigma=0.4$). This results in a more noise-robust set of linear measurements ($\text{OLM}_{\text{noise}}$), accompanied by a modest performance loss in low-noise regimes. The noise-optimized OLMs are less smooth than those optimized without noise (Figure˜4H). A more thorough investigation of noise robustness is worthwhile, but will require the development of a reconstruction algorithm that operates on noisy measurements.
@@ 3.5 Optimized linear measurements for perceptual loss
As an example, we computed a set of OLMs with respect to structural similarity index (SSIM) (Wang et al., 2004), a widely used perceptual image quality metric. Figure˜5 shows results obtained using this set of optimal linear measurements. We obtained a noticeable improvement in mean SSIM for a range of $k$, and this improvement applies to the majority of the images in the test set. In Figure˜5B, four example images are shown to highlight the visual improvement obtained by optimizing for the SSIM objective. However, it is worth noting that optimizing for SSIM does result in a lower average PSNR compared to the PCs, suggesting a trade-off between optimizing for perceptual quality versus MSE reconstruction, at least with our methods.
Lastly, in Figure˜5C, the SSIM-optimized measurement vectors are visualized, which appear quite distinct from both the PCs and their MSE-optimized counterpart shown in Figure˜4B. We further confirmed this difference by analyzing the measurement subspace, and found that the subspace distance between OLMs optimized toward different objectives is comparable to their distance from the PCs.
@@ 4 Related Work
The main advantages of our method are that (1) the linear measurements are optimized directly with respect to the end objective of the reconstruction problem, and that (2) we incorporate a highly expressive learned prior to exploit the higher-order statistics of natural images. As was shown in Section˜2, our method outperforms both random measurements and PCA.
@@ Appendix A Tweedie / Miyasawa Proof
For completeness, we provide the proof of the core relationship underlying diffusion models (Miyasawa, 1961; Robbins, 1992; Efron, 2011; Raphan & Simoncelli, 2007). Consider the problem of estimating a natural image $x$ given a noise-corrupted measurement $y=x+\epsilon$, with $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$. The estimate $\hat{x}(y)$ is chosen to minimize the mean of the squared error $x-\hat{x}(y)||_{2}^{2}$ over all images $x$ and observations $y$ corrupted with known $\sigma^{2}$. In Bayesian terms, the solution to this denoising problem is the posterior mean:
$\hat{x}(y)=\int x\ p(x|y)\ dx=\int x\ \frac{p(y|x)p(x)}{\int p(y|x)p(x)dx}\ dx.$ (10)
Miyasawa (1961) showed that the optimal estimator has a direct relationship to the score function of $\nabla\log p(y)$. First, write $p(y)=\int p(y|x)p(x)dx$. Assuming a Gaussian PDF for $p(y|x)$ and taking the gradient with respect to $y$ gives:
$\nabla_{y}p(y)=\frac{1}{\sigma^{2}}\int(x-y)p(y|x)p(x)dx.$ (11)
@@
Thus, the residual of the optimal denoiser, $\hat{x}(y)-y$ is proportional to the gradient of $\log p(y)$. Note that $p(y)$ is equal to the signal prior $p(x)$ convolved with (blurred by) a Gaussian with variance $\sigma^{2}$. This quantity can be used to sample from the implicit prior $p(x)$, for example, by using annealed Langevin dynamics, where the noise $\sigma^{2}$ corresponds to different temperature levels Song & Ermon (2019); Bussi & Parrinello (2007); Bhattacharya & Waymire (2009). See Appendix˜D for a detailed description of our sampling procedure.
@@ Appendix B Principal Components Analysis
Denote by $X$ the data matrix where the columns are individual mean-subtracted samples $x_{i}$. The linear projection is defined as $\hat{x_{i}}=MM^{T}x_{i}$, where $M$ is a matrix with orthogonal columns. Plugging this expression into Equation˜5, we have the expected squared-error loss for linear projection:
[…]
@@
When the data is in fact Gaussian distributed, linear projection onto the first $k$ PCs does achieve the minimal error possible, as shown in the bivariate example in Fig.˜1A. For non-Gaussian distributions, nonlinear estimation can potentially improve performance.
@@ Appendix C Sampling Algorithm
6:  $s_{t}\leftarrow f(x_{t-1})-x_{t-1}$ $\triangleright$ Compute the score from the denoiser residual
[…]
@@
10:  $x_{t}\leftarrow x_{t-1}+hs_{t}+\gamma_{t}z_{t}$ $\triangleright$ Perform a partial denoiser step and add noise
@@ Appendix D Constrained Sampling Algorithm
[…]
@@ E.3 MMSE estimate
We construct the linear inverse estimate by averaging multiple samples obtained using Algorithm˜2. The individual samples are shown in Figure˜2 and Supplementary Figure˜3 and Figure˜4. Averaging multiple samples achieves a lower MSE by approximating the posterior mean, which we use when reporting performance (PSNR values). The number of samples $n$ is set to $n=2$ when optimizing the linear measurement to optimize performance, as we found empirically that using more samples for the MMSE estimates during optimization does not improve performance during testing. For reporting performance on test data, we used $n=16$.
@@ E.4 Optimizing linear measurement
We search for the optimal linear measurement by performing stochastic gradient descent in the space of orthogonal matrices as described in the main text. The linear inverse procedure Algorithm˜2 is end-to-end differentiable, and thus optimization of the measurement matrix can be done directly in PyTorch by taking derivatives of reconstruction loss with respect to the matrix parameterization. In all cases, we used the Adam optimizer (Kingma & Ba, 2014) with a learning rate of $10^{-4}$ with exponential decay of $0.90$. We used a batch size of $64$, and training was run for $16$ epochs. The optimization was done on a single node in a GPU cluster, with 4 NVIDIA A100 GPUs. The optimization for a single OLM requires a few hours for the MNIST dataset, and about 24 hours for the CelebA dataset. The computational cost is primarily due to backpropagation through the diffusion model sampling procedure, which involves iterative applications of the U-Net denoiser.
@@ E.5 Subspace distance
[…]
$\text{subject to}\ \ u^{T}_{k}u_{j}=0,v^{T}_{k}v_{j}=0,\ \forall j=1,2,.,k-1.$
@@ E.6 Sequential OLM
The algorithm presented in Section˜2.4 learns an OLM independently for each value of $k$. Empirically, we observe that the linear measurement subspace of smaller $k$ is typically not fully contained in that of larger $k$. In this section, we modify the algorithm into a greedy version, where measurement vectors are added one at a time, each conditioned on the previous measurement matrix. This results in a sequentially ordered set of OLMs, analogous to the structure of PCA.
The algorithm works as follows: Given a measurement matrix $M_{k}\in R^{d\times k}$, we find a set of vectors $U=\{u_{1},u_{2},...,u_{d-k}\}$ as a basis of the null space of $M_{k}$. The new measurement vector $m^{*}$ is parameterized as $m^{*}=U\lambda$, where $\lambda$ is a unit vector. The next measurement matrix $M^{*}\in R^{d\times(k+1)}$ is obtained by concatenating $M_{k}$ with $m^{*}$ as $M^{*}=[M_{k}\ m^{*}]$. The objective function in Equation˜8 can be written with respect to $\lambda$ as:
[…]